A and b in a ring with 9 persons.ii the arrangement of 11 persons is at random then the probability that there are exactly 4 persons between a and b

Grade:12th pass

1 Answers

ANKIT
6 years ago
total no of ways to arrange 11 people in ring is 10!,   
now according to condition , let first arrange the 9 people by 8! ways , now for first person(assum a ) therir is 9 place to sheet , and for second person b there is only 2 place   ,( it can be done by two way by adjustion a,b ) so finaly require probability will 
(9*2*2*8!)/10! =4/10=2/5,  check it ,is it correct or not ?